OpenStudy (anonymous):

please help ASAP! part a. maria rented a coat at $285 for 3 days. if she rents the same coat for 6 days she has to pay $510. write an equation in standard form to represent the total rent (y) that Maria has to pay for renting the coat for x days. part b. write the equation obtained in part a using function notation part c. describe the steps to graph the equation obtained above on the coordinate axes. mention the label on the axes and the intervals

OpenStudy (anonymous):

Let (3,285) and (6,510) reprsent the coordinate points, where x = number of days coat is rented for; and y = price of rental Given these 2 points, we could write the equation of the straight line. We need the slope. Slope = (510-285)/(6-3) = 225/3 = 75 So the slope is 75. So the equation of the line is y - 285 = 75(x - 3) y - 285 = 75x - 225 y = 75x + 60 ---------------------------------------------------

OpenStudy (anonymous):

For part (b), same equation using function notation: f(x) = 75x + 60 would be in function notation ----------------------------------------------

OpenStudy (anonymous):

part (c) To graph this line...y = 75x + 60... Find the y -intercept...which is 60 (y = mx+b, b = y-intercept) Take another point, say x = 1, y = 75(1) + 60 = 135 So plot these 2 points, and you are done. I will graph it for you.

OpenStudy (anonymous):



OpenStudy (anonymous):

thank you so much I've been stuck on that problem for a while

OpenStudy (anonymous):

Look at the graph. I plotted the y-intercept of 60, and the point (1, 135). Connect the line. Notice that the line is not graphed for negative values of x, for in this problem x cannot be a negative number since it represents the number of days you rent the coat. And you cannot rent a coat for a negative number of days.

OpenStudy (anonymous):

Look all parts over, and make certain that all parts make 100% sense.

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

There is one adjustment to the graph I want to make. I would circle the point (0,60) on the graph as to not include that point. Becuase x = 0 means you are not renting the coat for any number of days. So why should y = 60..why pay 60 dollars for no rental. So exclude that point. And start after that point.

OpenStudy (anonymous):

I will draw the graph again.

OpenStudy (anonymous):



OpenStudy (anonymous):

Use the above graph. It shows rental starting with 1 day. Makes more practical sense than starting with 0 days. I used the point x = 2, y = 210. Use the equation y = 75x + 60. When x = 1, y = 135; when x = 2, y = 210. So those 2 points I plotted and connected the line.

OpenStudy (anonymous):

Do you follow???????